# Does Privacy Always Harm Fairness? Data-Dependent Trade-offs via Chernoff Information Neural Estimation

### v1

@@ Abstract
Fairness and privacy are two vital pillars of trustworthy machine learning. Despite extensive research on these individual topics, the relationship between fairness and privacy has received significantly less attention. In this paper, we utilize the information-theoretic measure Chernoff Information to highlight the data-dependent nature of the relationship among the triad of fairness, privacy, and accuracy. We first define Noisy Chernoff Difference, a tool that allows us to analyze the relationship among the triad simultaneously. We then show that for synthetic data, this value behaves in 3 distinct ways (depending on the distribution of the data). We highlight the data distributions involved in these cases and explore their fairness and privacy implications. Additionally, we show that Noisy Chernoff Difference acts as a proxy for the steepness of the fairness-accuracy curves. Finally, we propose a method for estimating Chernoff Information on data from unknown distributions and utilize this framework to examine the triad dynamic on real datasets. This work builds towards a unified understanding of the fairness-privacy-accuracy relationship and highlights its data-dependent nature.
@@ 1 Introduction
Yet while fairness and privacy have been deeply studied in isolation, their interaction remains less understood. Can models satisfy both simultaneously, or does enforcing one inevitably compromise the other? Recent work points to both incompatibility, exposing trade-offs between the two (Bagdasaryan et al., 2019; Chang and Shokri, 2021; Sanyal et al., 2022), and potential compatibility under specific assumptions (Cummings et al., 2019; Mangold et al., 2023).
In this work, we show that the answer ultimately lies in the structure of the data itself. We introduce the Noisy Chernoff Difference, a metric grounded in Chernoff Information, that captures the triad of fairness, privacy, and accuracy by quantifying how separability between groups evolves under privacy-preserving noise. Under Gaussian assumptions, we prove that fairness–privacy interactions fall into three distinct regimes: privacy may erode fairness, leave it unaffected, or even improve it, effectively granting “fairness for free” when distributional conditions align. To bridge theory with practice, we develop, to our knowledge, the first algorithm to estimate Chernoff Information directly from data, leveraging advances in density ratio estimation via neural networks (Choi et al., 2022), named Chernoff Information Neural Estimation (CINE in Section 5).
Beyond the Gaussian case, we validate our framework on three datasets, MNIST (LeCun et al., 1998), UCI Adult Income (Becker and Kohavi, 1996), and HSLS (Ingels et al., 2011), showing that the same distribution-driven patterns hold in practice. Our findings provide the first concrete evidence that whether fairness and privacy are in conflict or in harmony is not a universal law, but a property of the underlying data distribution. This reframes the fairness–privacy dilemma: it is not simply about balancing competing constraints, but about understanding when data itself enables both goals to be achieved in tandem, transforming how we think about building more prudent, trustworthy machine learning systems.
@@ 2 Related Works
Privacy and Fairness.
The relationship between fairness and privacy has attracted increasing attention in recent years. Bagdasaryan et al. (2019) show that differential privacy disproportionately harms the accuracy of underrepresented groups, while Chang and Shokri (2021) demonstrate that minority groups are more vulnerable to membership inference attacks. Cummings et al. (2019) prove the impossibility of achieving exact fairness and DP with nontrivial accuracy, but propose an algorithm that satisfies DP with approximate fairness under relaxed constraints. Sanyal et al. (2022) show that under long-tailed, imbalanced distributions and strict privacy, enforcing both fairness and privacy worsens accuracy, whereas Mangold et al. (2023) establish that fairness disparities between private and non-private models decay inversely with sample size. Several works also develop algorithms jointly ensuring fairness, privacy, and accuracy (Lyu et al., 2020; Lowy et al., 2023; Ghoukasian and Asoodeh, 2024; Jagielski et al., 2019). For a comprehensive survey, see Yao and Juarez (2025).
@@
Chernoff Information, introduced by Chernoff (1952), is closely tied to Bayes error bounds. Nielsen (2011, 2013, 2022) provides key properties we exploit throughout this work. Although information-theoretic tools have been widely applied to fairness and privacy (Ghassami et al., 2018), Chernoff Information itself has rarely been used in these domains. Dutta et al. (2020) show that disparities in Chernoff Information imply a trade-off between equal opportunity fairness and accuracy, while Ünsal and Önen (2024) define Chernoff Differential Privacy, where output distributions for neighboring inputs must have bounded Chernoff Information. To our knowledge, ours is the first work to connect Chernoff Information with the joint analysis of accuracy, fairness, and privacy, and to provide an estimator beyond synthetic settings.

### v2

@@ Abstract
Fairness and privacy are two vital pillars of trustworthy machine learning. Despite extensive research on these individual topics, their relationship has received significantly less attention. In this paper, we utilize an information-theoretic measure Chernoff Information to characterize the fundamental trade-off between fairness, privacy, and accuracy, as induced by the input data distribution. We first propose Chernoff Difference, a notion of data fairness, along with its noisy variant, Noisy Chernoff Difference, which allows us to analyze both fairness and privacy simultaneously. Through simple Gaussian examples, we show that Noisy Chernoff Difference exhibits three qualitatively distinct behaviors depending on the underlying data distribution. To extend this analysis beyond synthetic settings, we develop the Chernoff Information Neural Estimator (CINE), the first neural network–based estimator of Chernoff Information for unknown distributions. We apply CINE to analyze the Noisy Chernoff Difference on real-world datasets. Together, this work fills a critical gap in the literature by providing a principled, data-dependent characterization of the fairness–privacy interaction.
@@ 1 Introduction
Figure 1: Summary of our contributions. Starting from group-specific class-conditional distributions (left), we use CINE to estimate their Chernoff Information (middle), and compute the Chernoff Difference to diagnose whether privacy can help or hurt fairness (right).
Yet while fairness and privacy have been deeply studied in isolation, their interaction remains less understood. Can models satisfy both simultaneously, or does enforcing one inevitably compromise the other? Recent work points to both incompatibility, exposing trade-offs between the two (Bagdasaryan et al., 2019; Chang and Shokri, 2021; Tran et al., 2021b; Sanyal et al., 2022), and potential compatibility under specific assumptions (Cummings et al., 2019; Mangold et al., 2023; Shaham et al., 2025). In this work, we show that the answer can lie in the structure of the data itself. We make the following contributions:
  • We introduce Chernoff Difference (CD), a metric grounded in Chernoff Information, which serves as a notion of data fairness by characterizing how groups differ in their classification separability. We then propose Noisy Chernoff Difference, an extension of CD that captures the fairness–privacy–accuracy triad by modeling how group separability evolves under privacy-preserving noise.
  • Under Gaussian assumptions, we prove that fairness–privacy interactions fall into three distinct regimes: privacy may erode fairness, leave it unaffected, or even improve it, effectively granting “fairness for free” when distributional conditions align.
  • To bridge theory with practice, we develop, to our knowledge, the first algorithm to estimate Chernoff Information directly from data, leveraging advances in density ratio estimation via neural networks (Choi et al., 2022), named Chernoff Information Neural Estimation (CINE in Section 6).
  • We apply CINE on more complex datasets, mixture of Gaussians, MNIST (LeCun et al., 1998), and UCI Adult (Becker and Kohavi, 1996), and demonstrate that CD can capture the fairness–privacy dynamics in a wide range of settings.
To the best of our knowledge, our findings provide the first concrete evidence that whether fairness and privacy are in conflict or in harmony is not a universal law, but a property of the underlying data distribution, thereby building the core foundation of a data-dependent understanding of the complex relationships between fairness, privacy, and accuracy.
@@ 2 Related Works
Privacy and Fairness. The relationship between fairness and privacy has attracted increasing attention in recent years. Bagdasaryan et al. (2019) show that DP disproportionately harms the accuracy of underrepresented groups, while Chang and Shokri (2021) demonstrate that minority groups are more vulnerable to membership inference attacks. Tran et al. (2021a) demonstrate that privacy can amplify disparities for groups closer to the decision boundary. Cummings et al. (2019) prove the impossibility of achieving exact fairness and DP with nontrivial accuracy, but propose an algorithm that satisfies DP with approximate fairness under relaxed constraints. Sanyal et al. (2022) show that under long-tailed, imbalanced distributions and strict privacy, enforcing both fairness and privacy worsens accuracy, whereas Mangold et al. (2023) establish that fairness disparities between private and non-private models decay inversely with sample size. Several works also develop algorithms jointly ensuring fairness, privacy, and accuracy (Lyu et al., 2020; Lowy et al., 2023; Ghoukasian and Asoodeh, 2024; Jagielski et al., 2019). For a survey, see Fioretto et al. (2022) and Shaham et al. (2025). Despite these advances, there is still a limited understanding of how the input data distribution shapes this relationship, as highlighted in a recent review paper (Yao and Juarez, 2025). Our work addresses this gap by providing a distribution-level characterization of how privacy reshapes group separability, offering a principled explanation of when privacy may harm or help fairness.
@@
Chernoff Information, introduced by Chernoff (1952), determines the optimal error exponent of Bayes-optimal classification. Nielsen (2011, 2013, 2022) provides key properties we exploit throughout this work. Although information-theoretic tools have been applied to fairness and privacy (Ghassami et al., 2018), Chernoff Information itself has rarely been used in these domains. Dutta et al. (2020) show that disparities in Chernoff Information imply a trade-off between equal opportunity fairness and accuracy, while Ünsal and Önen (2024) define Chernoff DP, where output distributions for neighboring inputs must have bounded Chernoff Information. To our knowledge, our work is the first to connect Chernoff Information with the joint analysis of accuracy, fairness, and privacy, and to provide an estimator beyond synthetic settings.